OpenStudy (anonymous):

if f(x)=3x+1 and g(x)= 2/3x-2 find f(g)(9))

OpenStudy (anonymous):

f(g(9)) can u tell me wat g(9) would be?

OpenStudy (anonymous):

@Saikam it literally says find f(g(9))

OpenStudy (anonymous):

I was giving u a hint. first find g(9) then find f(this value) want me to solve it for u?

OpenStudy (anonymous):

yes please!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

g(9) = 2/(3*9 - 2) = 2/25 so f(g(9)) = f(2/25) = 3(2/25) + 1 6/25 + 1 (6+25) / 25 = 31/25

OpenStudy (anonymous):

thankyou @Saikam !!!!!!!!!!!!!!!